# Network Oblivious Transfer via Noisy Channels: Limits and Capacities

### v2

@@ Appendix E Proof of Theorem 10
$H(M_{i\overline{Z^{\prime}}_{i}}|V,Z_{i},Z^{\prime}_{i},M_{iZ^{\prime}_{i}})=H(M_{i\overline{Z^{\prime}}_{i}}).$ (47)
This implies that given the information $V$, the choice $Z_{i}$, the effective choice $Z^{\prime}_{i}$ and $M_{iZ^{\prime}_{i}}$, Bob has no additional information about $M_{i\overline{Z^{\prime}}_{i}}$. Consider the mutual information between Alice’s bits and the information received by Bob $I(X_{i};V)=I(M_{i0},M_{i1};V)$. Since Bob should learn $M_{iZ^{\prime}_{i}}$ and nothing about $M_{i\overline{Z^{\prime}}_{i}}$, we have:
$I(M_{iZ^{\prime}_{i}}$ $,M_{i\overline{Z^{\prime}}_{i}};V|Z_{i},Z^{\prime}_{i},M_{iZ^{\prime}_{i}})=I(M_{iZ^{\prime}_{i}};V|Z_{i},Z^{\prime}_{i},M_{iZ^{\prime}_{i}})+I(M_{i\overline{Z^{\prime}}_{i}};V|Z_{i},Z^{\prime}_{i},M_{iZ^{\prime}_{i}}).$ (48)
The first term should not equal zero because $V$ must also convey the value of $M_{iZ^{\prime}_{i}}$ to Bob. The second term should be zero due to the perfect secrecy of OT. So consider the second term:
$I(M_{i\overline{Z^{\prime}}_{i}};V|Z_{i},Z^{\prime}_{i},M_{iZ^{\prime}_{i}})$ $=I(M_{i\overline{Z^{\prime}}_{i}};V|Z_{i},Z^{\prime}_{i})$
$=H(M_{i\overline{Z^{\prime}}_{i}}|Z_{i},Z^{\prime}_{i})-H(M_{i\overline{Z^{\prime}}_{i}}|V,Z_{i},Z^{\prime}_{i})$
$=H(M_{i\overline{Z^{\prime}}_{i}}|Z_{i},Z^{\prime}_{i})-H(M_{i\overline{Z^{\prime}}_{i}})$
$=-I(M_{i\overline{Z^{\prime}}_{i}};Z_{i},Z^{\prime}_{i}),$ (49)
where the first equality is due the fact that $M_{i\overline{Z^{\prime}}_{i}}$ and $M_{iZ^{\prime}_{i}}$ are independent, and the third equality follows from (47).
As $I(M_{i\overline{Z^{\prime}}_{i}};Z_{i},Z^{\prime}_{i})$ is a non-negative quantity, then we proved that $I(M_{i\overline{Z^{\prime}}_{i}};Y|M_{iZ^{\prime}_{i}},Z_{i},Z^{\prime}_{i})$ is not a positive quantity, unless $I(M_{i\overline{Z^{\prime}}_{i}};Z_{i},Z^{\prime}_{i})=0$, which implies that all inputs $(X_{i},Z_{i})$ are independent. Combining (48) and (E), we have:
$I(M_{iZ^{\prime}_{i}},M_{i\overline{Z^{\prime}}_{i}};V|Z_{i},Z^{\prime}_{i},M_{iZ^{\prime}_{i}})$ $=I(M_{iZ^{\prime}_{i}};V|Z_{i},Z^{\prime}_{i})$
$=H(M_{iZ^{\prime}_{i}}|Z_{i},Z^{\prime}_{i})-H(M_{iZ^{\prime}_{i}}|V,Z_{i},Z^{\prime}_{i})$
$=H(M_{iZ^{\prime}_{i}}|Z_{i},Z^{\prime}_{i}),$ (50)
where the second equality is because given $V$ and $Z^{\prime}_{i}$, the uncertainty about $M_{iZ^{\prime}_{i}}$ is zero. The term $H(M_{iZ^{\prime}_{i}}|Z_{i},Z^{\prime}_{i})$ is not equal to zero because we already assumed that $M_{iZ^{\prime}_{i}}$ and $Z^{\prime}_{i}$ are independent. This is a contradiction. ∎
@@ Appendix H Proof of Corollary 14
for $i\in\{1,2\}$, where $(a)$ follows from the security criteria, $(b)$ follows from Fano’s lemma, $(c)$ and $(d)$ follow from the fact that $C_{i}=f_{i}(S_{i},E^{n})$, $(e)$ follows from the Markov chain $Y^{n}-(X_{1}^{n},X_{2}^{n})-(S_{1},S_{2})$, and $(f)$ follows from the memoryless property of the channel and an argument similar to [42, Theorem 4]. As is mentioned before and proved in [42] for the case of constant random variable $E$, we can remove the impact of $E$ from the above mutual information quantity. The whole above process can similarly be repeated for the joint entropy $H(S_{1},S_{2})$. This completes the proof. ∎
@@ I-B The second upper bound on OT capacity
    where $(a)$ follows from (73), $(b)$ is because conditioning does not increase the entropy, and $(c)$ follows from the arguments presented in [5]. Similarly, we can prove that $k_{1}+k_{2}=H(M_{1\overline{z}_{1}},M_{2\overline{z}_{2}}|Z_{1}=z_{1},Z_{2}=z_{2})\leq\sum_{l=1}^{n}H(X_{1,l},X_{2,l}|Y_{l})+o(n)$.

### v4

@@ Appendix E Proof of Theorem 10
$H(M_{i\overline{Z^{{}^{\prime}}_{i}}}|V,Z_{i},Z^{\prime}_{i},M_{iZ^{{}^{\prime}}_{i}})=H(M_{i\overline{Z^{{}^{\prime}}_{i}}}).$ (51)
This implies that given the information $V$, the choice $Z_{i}$, the effective choice $Z^{\prime}_{i}$ and $M_{iZ^{{}^{\prime}}_{i}}$, Bob has no additional information about $M_{i\overline{Z^{{}^{\prime}}_{i}}}$. Consider the mutual information between Alice’s bits and the information received by Bob $I(X_{i};V)=I(M_{i0},M_{i1};V)$. Since Bob should learn $M_{iZ^{{}^{\prime}}_{i}}$ and nothing about $M_{i\overline{Z^{{}^{\prime}}_{i}}}$, we have:
$I(M_{iZ^{{}^{\prime}}_{i}}$ $,M_{i\overline{Z^{{}^{\prime}}_{i}}};V|Z_{i},Z^{\prime}_{i},M_{iZ^{{}^{\prime}}_{i}})=I(M_{iZ^{{}^{\prime}}_{i}};V|Z_{i},Z^{\prime}_{i},M_{iZ^{{}^{\prime}}_{i}})+I(M_{i\overline{Z^{{}^{\prime}}_{i}}};V|Z_{i},Z^{\prime}_{i},M_{iZ^{{}^{\prime}}_{i}}).$ (52)
The first term should not equal zero because $V$ must also convey the value of $M_{iZ^{{}^{\prime}}_{i}}$ to Bob. The second term should be zero due to the perfect secrecy of OT. So consider the second term:
$I(M_{i\overline{Z^{{}^{\prime}}_{i}}};V|Z_{i},Z^{\prime}_{i},M_{iZ^{{}^{\prime}}_{i}})$ $=I(M_{i\overline{Z^{{}^{\prime}}_{i}}};V|Z_{i},Z^{\prime}_{i})$
$=H(M_{i\overline{Z^{{}^{\prime}}_{i}}}|Z_{i},Z^{\prime}_{i})-H(M_{i\overline{Z^{{}^{\prime}}_{i}}}|V,Z_{i},Z^{\prime}_{i})$
$=H(M_{i\overline{Z^{{}^{\prime}}_{i}}}|Z_{i},Z^{\prime}_{i})-H(M_{i\overline{Z^{{}^{\prime}}_{i}}})$
$=-I(M_{i\overline{Z^{{}^{\prime}}_{i}}};Z_{i},Z^{\prime}_{i}),$ (53)
where the first equality is due the fact that $M_{i\overline{Z^{{}^{\prime}}_{i}}}$ and $M_{iZ^{{}^{\prime}}_{i}}$ are independent, and the third equality follows from (51).
As $I(M_{i\overline{Z^{{}^{\prime}}_{i}}};Z_{i},Z^{\prime}_{i})$ is a non-negative quantity, then we proved that $I(M_{i\overline{Z^{{}^{\prime}}_{i}}};Y|M_{iZ^{{}^{\prime}}_{i}},Z_{i},Z^{\prime}_{i})$ is not a positive quantity, unless $I(M_{i\overline{Z^{{}^{\prime}}_{i}}};Z_{i},Z^{\prime}_{i})=0$, which implies that all inputs $(X_{i},Z_{i})$ are independent. Combining (52) and (E), we have:
$I(X_{i};V|Z_{i},Z^{\prime}_{i},M_{iZ^{{}^{\prime}}_{i}})$ $\geq I(M_{iZ^{{}^{\prime}}_{i}},M_{i\overline{Z^{{}^{\prime}}_{i}}};V|Z_{i},Z^{\prime}_{i},M_{iZ^{{}^{\prime}}_{i}})$
$=I(M_{iZ^{{}^{\prime}}_{i}};V|Z_{i},Z^{\prime}_{i})$
$=0,$ (54)
where the first equality is because given $V$ and $Z^{\prime}_{i}$, the uncertainty about $M_{iZ^{{}^{\prime}}_{i}}$ is zero. ∎
@@ Appendix H Proof of Corollary 14
for $i\in\{1,2\}$, where $(a)$ follows from the security criteria, $(b)$ follows from Fano’s lemma, $(c)$ and $(d)$ follow from the fact that $C_{i}=f_{i}(S_{i},E^{n})$, $(e)$ follows from the Markov chain $Y^{n}-(X_{1}^{n},X_{2}^{n})-(S_{1},S_{2})$, and $(f)$ follows from the memoryless property of the channel and an argument similar to [Maurer, Theorem 4]. As is mentioned before and proved in [Maurer] for the case of constant random variable $E$, we can remove the impact of $E$ from the above mutual information quantity. The whole above process can similarly be repeated for the joint entropy $H(S_{1},S_{2})$. This completes the proof. ∎
@@ I-B The second upper bound on OT capacity
    where $(a)$ follows from (77), $(b)$ is because conditioning does not increase the entropy, and $(c)$ follows from the arguments presented in [Rudolf1]. Similarly, we can prove that $k_{1}+k_{2}=H(M_{1\overline{z}_{1}},M_{2\overline{z}_{2}}|Z_{1}=z_{1},Z_{2}=z_{2})\leq\sum_{l=1}^{n}H(X_{1,l},X_{2,l}|Y_{l})+o(n)$.